OpenStudy (anonymous):

the integral of 2^sqrt(x)?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\int 2^{\sqrt{x}}dx\]?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

then it is parts you need , integration by parts

OpenStudy (anonymous):

sorry for wrong interpretation,put\[\sqrt{x}=t,x=t^2,dx=2tdt\] \[I=\int\limits 2^\sqrt{x}dx\] \[=\int\limits 2^t(2t)dt=\int\limits (2t)2^t dt\] integrate by parts ?

OpenStudy (misty1212):

\[u=\sqrt{x},, du=\frac{1}{2}\sqrt{x}dx\] and the integral becomes \[2\int 2^udu\]

OpenStudy (misty1212):

oops i meant \[2\int 2^uudu\]

OpenStudy (misty1212):

or what @surjithayer wrote

OpenStudy (anonymous):

Thank You for your answers